Solve the following differential equation: 2dy/dx + y = y^3(x - 1)

1
Expert's answer
2022-05-11T17:25:10-0400

"y'+y\/2=(xy^3)\/2-y^3\/2"

Bernuli equalation

"y'+a(x)y=b(x)y^n"

a(x)=1/2

b(x)=(x-1)/2

n=3


"y'\/y^3+1\/(2y^2)=x\/2-1\/2"

Deleting, we miss solution y=0


"u=1\/y^2"

"u'=-2y'\/y^3"

"y=1/\sqrt u"

"y'=-u'y^3\/2"

"u\/2-u'\/2=x\/2-1\/2"

u-u'=x-1

-u'=x-u-1


v=x-u-1

v'=1-u'

u=x-v-1

u'=1-v'

v'-1=v

v'=v+1=dv/dx

dv=(v+1)dx

dv/(v+1)=dx

Deleting, we miss solution v+1=0, x=1/y2

"\int dv/(v+1)=\int dx"

ln (v+1)=x+C

v+1=ex+C

x-u=Cex

x-1/y2=Cex

"y=\sqrt {\frac{1}{x-Ce^x}}"